If someone could please help with the following. I am lost where to start.
Two cars leave an intersection. One car travels north, the other east. When the car travelling north had gone 9km, the distance between the cars was 6km more than the distance travelled by the car heading east. How far had the eastbound car travelled.
Thank you.

Answer by edjones(7311)

We have a right triangle.
Let a=9, c=b+6
a^2+b^2=c^2
9^2+b^2=(b+6)^2
9^2+b^2=b^2+12b+36
12b-45=0
3(4b-15)=0
4b=15
b=3.75 km
.
Check:
9^2+3.75^2=9.75^2
81+14.0625=95.0625 true
.
Ed